# Modeling Human Behavior in a Strategic Network Game with Complex Group Dynamics

### v2

@@ Abstract
Human networks greatly impact important societal outcomes, including wealth and health inequality, poverty, and bullying. As such, understanding human networks is critical to learning how to promote favorable societal outcomes. As a step toward better understanding human networks, we compare and contrast several methods for learning, from a small data set, models of human behavior in a strategic network game called the Junior High Game (JHG) [41]. These modeling methods differ with respect to the assumptions they use to parameterize human behavior (behavior vs. community-aware behavior) and the moments they model (mean vs. distribution). Results show that the highest-performing method, called hCAB, models the distribution of human behavior rather than the mean and assumes humans use community-aware behavior rather than behavior matching. When applied to small societies (6-11 individuals), the hCAB model closely mirrors the population dynamics of human groups (with notable differences). Additionally, in a user study, human participants were unable to distinguish individual hCAB agents from other humans, thus illustrating that the hCAB model also produces plausible (individual) human behavior in this strategic network game.
1 Introduction
@@
Human networks, collections of people linked through social connections, influence how individuals exchange resources [26], form alliances [12, 20], and navigate competitive pressures [8, 46]. These networks, which are studied in many disciplines [43], exhibit emergent behaviors related to inequality [6, 11, 37], information flow [9] and social behaviors [40, 10, 45]. While the study of human networks has yielded deep insights [29, 13, 5, 35], understanding the precise decision-making processes that drive human behavior within these networks remains a major challenge [49]. What determines whether an individual cooperates or competes? How do local interactions scale up to produce global patterns of stability, inequality, or conflict? How might an individual change their behavior to encourage preferred social change? Addressing these questions requires computational models that capture the complexities of human decision-making in networked environments.
@@
In human networks, each individual makes decisions based on personal incentives, past experiences, and the social landscape. Sometimes individual incentives align with the collective good, like the invisible hand [42], but other times they are diametrically opposed, like the free rider problem [34]. These individual actions lead to emergent behaviors that shape societal outcomes. Extensive studies have explored emergent properties of human networks [22, 23, 39], investigating, for example, how randomness affects network structures [14], how power distributes among individuals [6], and how self-organization arises within different social [47] and economic settings [1, 48]. These macroscopic perspectives on how human societies function are valuable, yet they often are insufficient for fully modeling human networks and how they evolve. More detailed models of human decision-making are needed to better understand human networks.
To help address this challenge, we consider modeling human behavior in the Junior High Game (JHG) [41], a controlled game environment designed to capture the strategic trade-offs in human networks. In this game, players engage in repeated economic exchanges, choosing whether to give, keep, or take resources from others. Decisions are shaped by social factors such as dyadic relationships [17, 15] triadic (indirect) interactions [18, 33], and group considerations [31, 7, 21]. JHG scenarios produce networks that are dynamic, directed, weighted, and signed, thus providing a rich testbed for studying human behavior in networked settings.
@@
A fundamental challenge in modeling human decision-making in the JHG lies in determining how to parameterize and represent behavior [36]. In this paper, we consider models that make different assumptions about human decision-making. One approach, behavior matching [4], assumes that agents imitate the observed actions of others, while another approach, community-aware behavior [41], considers broader structural and social cues within the network. We also examine how modeling different statistical moments (the mean verses the distribution of behavior) impact model performance. Together, this produces four different algorithms for modeling human behavior in the JHG, which we compare and contrast in small-scale societies (6-11 players).
We evaluate the ability of the resulting modeling algorithms to produce human behavior in the JHG in two ways. First, we compare and contrast the population dynamics of agent societies with those of human groups. Results show that societies of one of these agents, called hCAB, have population dynamics that closely mirror (with some differences) human populations. Second, we conduct a user study to identify the extent to which individual hCAB agents behave like human players. Results of this study show that participants were unable to distinguish hCAB agents from other human players. These findings not only inform the development of models that can simulate human behavior in the JHG, but also offer perspectives on the mechanisms that drive the dynamics of real-world networks.
2 A strategic network game to model human networks
In this paper, we model human networks using the Junior High Game (JHG) [41]. The JHG models key properties of human networks including conflicting interests, a large and flexible strategy space, high interdependence among agents, power asymmetry, and (often) non-convergence [23, 44]. The JHG induces a dynamic (time varying), directed, weighted, and signed network.
We paraphrase Skaggs et al. [41] to review this game: The JHG is played by a set of players $I$ over a series of rounds in which each player $j\in I$ seeks to become popular. Initially, every player is assigned a popularity, which changes over time based on interactions, represented by token allocations, among the players. In every round, each player $j$ allocates $N$ tokens, each of which it can keep, give to another player, or attack another player. Keeping tokens positively impacts $j$’s subsequent popularity, while tokens $j$ gives to player $i$ positively impact $i$’s subsequent popularity (but do nothing for $j$). Finally, when $j$ attacks $i$, it negatively impacts $i$’s subsequent popularity while positively impacting $j$’s. After all players allocate their tokens in a round, the players’ popularities are updated based on the players’ token allocations. A new round then begins.
The amount that a token transaction impacts subsequent popularity depends on the popularity of the player allocating the token. For example, receiving a token from a more popular individual results in a higher increase in popularity than receiving a token from a less popular individual. Thus, players with higher popularity can have higher impact than those with lower popularity.
(a)                                            (b)                                 (c)
@@ 2 A strategic network game to model human networks
Figure 1: Example scenario illustrating various JHG dynamics. (a) Popularity of the players over time. (b) Token allocations in rounds 1-4. Red dashed arrows indicate tokens used to attack while black solid arrows indicate tokens given. (c) Token allocations in rounds 5-6.
@@
To increase understanding of JHG dynamics, consider the contrived example depicted in Figure 1. First, when players exchange (give) tokens with each other, they rise in popularity (e.g., players p2 and p3). Second, when a player gives tokens without receiving as much in return, they tend to decrease in popularity (e.g., p4). Third, when a player keeps its token, it maintains much of its popularity, but does not increase in popularity in the absence of other interactions (e.g., p1 in rounds 1-4). Fourth, attacking those not keeping tokens results in the attacker becoming more popular and the victim becoming less popularity (e.g., p5 attacking p7 and p8 in rounds 1-4). When other players do not keep or retaliate, attacking is the fastest way to gain popularity. It also increases the attacker’s relative popularity in relation to the victim. Fifth, keeping tokens blocks attacks. For example, p6’s attacks on p1 in rounds 1-4 are blocked because p1 is keeping its tokens–p1 experiences no loss, while p6 quickly loses popularity due to not gaining from its token allocations. Finally, attacks that are stronger than the defense (strength is measured as tokens allocated multiplied by player popularity) result in inflicted damage. For example, p5’s attack, combined with p6’s, on p1 in rounds 5-6 is substantially stronger than p1’s defense. Thus, p1 loses popularity and p5 and p6 profit somewhat.
@@
While the contrived example presented above deals primarily with pairwise interactions for illustrative purposes, these dynamics can give rise to complex group behavior and emergent social norms. Frequently, alliances form and dissolve as players join together to help each other, punish those that violate norms, mitigate powerful enemies, and exploit others.
3 Modeling human behavior in the JHG using human data
@@
In this section, we introduce four algorithms for learning models of human behavior in the JHG from a relatively small data set of games played by experienced human players. These algorithms vary based on assumptions encoded in their parameter-based representations and their methods for learning parameterizations. We discuss these design choices in turn.
3.1 Parameter-based models of behavior
We construct algorithms based on two parameter-based models of behavior for playing the JHG: Behavior Matching (TFT) and community-aware behavior (CAB).
@@ 3.1 Parameter-based models of behavior
Behavior Matching (TFT): In behavior matching, or tit-for-tat (TFT), a player reciprocates the actions that other players directed towards them in the previous round [4]. We consider this model due to its prevalence in prior work. Tit-for-tat and its variants have repeatedly been shown to be effective and robust in repeated prisoner’s dilemmas (e.g., [3, 32, 19]), a common domain for studying when and how humans and other agents cooperate with each other.
@@
Defining TFT in the JHG is not straightforward, as a player may receive more or less in a round than it has the ability to reciprocate. In such cases, an agent must determine who to exclude (when receiving too much) or what to do with extra tokens (when not receiving enough). We have defined a TFT algorithm for the JHG whose behavior is tuned with seven parameters. These parameters impact how the agent allocates tokens in the first round, what it seeks to match (token allocations or influence), and how its deals with excess and shortfalls. Algorithmic details are given in SM-B.
@@
Community-Aware Behavior (CAB): The second parameter-based model we consider is the CAB (Community Aware Behavior) algorithm [41]. While behavior matching assumes humans primarily reciprocate the actions of others, the CAB agent is based primarily on group dynamics. A CAB agent determines token allocations in a round using a three-step process. First, it assigns players to groups based on past token allocations. Second, it chooses a (potentially new) group it would like to form, support, or join. Finally, it selects token allocations designed to build up its selected group, establish itself in the group, and deal with out-of-group threats. CAB behavior is defined by thirty parameters which determine how it identifies existing groups, chooses a group to try to be part of, and allocates tokens to build its group. We use the CAB implementation provided by Skaggs et al. [41].
@@
Assuming that human behavior in the JHG can be modeled using either of these algorithms is most certainly incorrect at some level. However, identifying the extent to which these models can and cannot approximate human behavior can shed valuable insights into human networks.
3.2 Learning parameterizations
@@
To determine how well either parameter-based model (TFT or CAB) can model human behavior in the JHG, we learn parameterizations of each model from a limited amount of human data using two different methods: Particle Swarm Optimization (PSO) [25] and Evolutionary Population Distribution Modeling (EPDM), a methodology proposed in this paper. Both algorithms take as input a parameter-based model and output one or more parameterizations of the model. Let $m$ be the number of parameters in the desired model. Then, $\Theta\in(\theta_{1},\theta_{2},\cdots,\theta_{m})$ is a parameterization of the model in which $\theta_{i}$ is the value of the $i^{\rm th}$ parameter.
@@
Particle Swarm Optimization (PSO): In PSO, a set of particles (i.e., parameterizations) are first randomly generated. Over time, these parameterizations are adapted to both explore the parameter space and converge toward the best parameterization. In our case, the best parameterization is determined by minimizing the error, over all data samples, between the token allocations made by human players and the token allocations made by the model (given the parameterization specified by the particle) in the same situations. In this way, PSO produces a parameterization that describes mean behavior in the data set.
@@
We experimented with two different error (or distance) functions: (1) mean squared error and (2) a custom built function in which the error is based on how well token allocations match across a variety of different properties (defined in SM-D). Because the second function worked best in all cases, we used that error function in all of our experiments.
Evolutionary Population Distribution Modeling (EPDM): Because mean human behavior, when followed by all individuals, does not necessarily produce human-like population dynamics or realistic individual behavior, we introduce a second method, called EPDM, for modeling human behavior from data. EPDM attempts to model the distribution of human behavior, rather than mean behavior.
@@
The EPDM algorithm, outlined in Algorithm 1, receives as input a parameter-based behavior model $\mathcal{M}$ (TFT or CAB) and a set of player games $\Gamma$ (from the training data). Each player game $\gamma\in\Gamma$ specifies the token allocations made by a human player in each round $t$ of a game, along with the game state available to the player at the beginning of each round (specifically, ${\mathcal{P}}(t)$, ${\mathcal{I}}(t)$, and $X(t)$). The algorithm outputs a set of $N$ (we used 100) parameterizations of $\mathcal{M}$ that best estimate the distribution of player strategies in $\Gamma$.
Algorithm 1 EPDM – Computes $N$ parameterizations to estimate a distribution of strategies
@@
EPDM learns this distribution of parameterizations using genetic evolution. Initially, a set of $N$ parameterizations (denoted $\Pi(0)$) are randomly generated. Over a series of $G$ generations, the pool of parameterizations is evolved to better estimate the distribution of strategies used by human players in the supplied player games. Each generation $g$ proceeds in three steps. In step 1, the set of player games for which parameterization $\pi\in\Pi(g)$ is a top performer is computed. This set, denote $\Gamma^{\pi}(g)$, is determined as follows. Let $a^{\gamma}(t)$ denote the token allocations made in player game $\gamma\in\Gamma$ in round $t$. Also, let $a^{\pi}(t)$ be the allocations that parameterization $\pi$ would make in round $t$ given the history of game transactions made prior to round $t$. Then, the average error of $\pi$ in player game $\gamma$ is
$\delta^{\pi}_{\gamma}=\frac{1}{T_{\gamma}}\sum_{t=1}^{T_{\gamma}}dist(a^{\gamma}(t),a^{\pi}(t)),$ (1)
@@
where $T_{\gamma}$ is the number of rounds in player game $\gamma$ and $dist(a,b)$ is the error function specifying the distance between allocations $a$ and $b$ (see SM-D for details). $\pi$ is a top-performing parameterizations for player game $\gamma$ if $\delta^{\pi}_{\gamma}-\min_{\pi^{\prime}\in\Pi}\delta^{\pi^{\prime}}_{\gamma}<\varepsilon$. We used $\varepsilon=0.01$.

### v3

@@ Abstract.
Human networks greatly impact important societal outcomes, including wealth and health inequality, poverty, and bullying. As such, understanding human networks is critical to learning how to promote favorable societal outcomes. As a step toward better understanding human networks, we compare and contrast several methods for learning models of human behavior in a strategic network game called the Junior High Game (JHG) (skaggs2024fostering). These modeling methods differ with respect to the assumptions they use to parameterize human behavior (behavior matching vs. community-aware behavior) and the moments they model (mean vs. distribution). Results show that the highest-performing method, called hCAB, models the distribution of human behavior rather than the mean and assumes humans use community-aware behavior rather than behavior matching. When applied to small societies, the hCAB model closely mirrors the population dynamics of human groups (with notable differences). Additionally, in a user study, human participants had difficulty distinguishing hCAB agents from other humans, thus illustrating that the hCAB model also produces plausible (individual) behavior in this strategic network game.
Key words and phrases:
Complex systems, emergent behavior, multi-agent systems, human modeling, network science
doi:
NQKD4743
1. Introduction
@@
Human networks (collections of interconnected people) influence how individuals exchange resources willer1999network, form alliances (cranmer2012toward; gulati1995social), and navigate competitive pressures (burt2005brokerage; vu2006learning). These networks, which are studied in many disciplines (svendsen2009handbook; levin1998ecosystems), exhibit emergent behaviors related to inequality (barabasi1999emergence; salganik2006experimental), information flow (Harush2017), and social behaviors (shoham1997emergence; christakis2010social; centola2010spread; centola2018behavior; veenstra2021social). While the study of human networks has yielded deep insights (newman2003structure; easley2010networks; bar2019dynamics; pagan2019game), understanding the precise decision-making processes that drive human behavior within these networks remains a major challenge (young2015evolution). What determines whether an individual cooperates or competes? How do local interactions scale up to produce global patterns of stability, inequality, or conflict? How might an individual change their behavior to encourage preferred social change? Addressing these questions requires computational models that capture the complexities of human decision-making in networked environments.
@@
In human networks, individuals make decisions based on personal incentives, past experiences, and the social landscape. Sometimes individual incentives align with the collective good, like the invisible hand (smith1776wealth), but other times they are diametrically opposed, as in the free-rider problem (olson1965logic). These individual actions lead to emergent behaviors that shape societal outcomes. Extensive studies have explored emergent properties of human networks (holland1995hidden; jensen2022complexity; shoham2008multiagent), investigating, for example, how randomness affects network structures (erdos1959random), how power distributes among individuals (barabasi1999emergence), and how self-organization arises within different social (watts1998collective) and economic settings (arthur1994inductive; young2001individual). These macroscopic perspectives on how human societies function are valuable, yet they are often insufficient for fully understanding human networks and how they evolve. More detailed models of human decision-making on networks are needed.
To help address this challenge, we model human behavior in a strategic network game designed to capture the emergent properties of complex (networked) societies, including wealth and power distributions and mixing patterns defined by dyadic (gouldner1960norm; fehr2000fairness), triadic (granovetter1973strength; nowak2005evolution), and group dynamics (nowak2006five; bowles2011cooperative; henrich2015secret). These dynamics emerge due to power asymmetry, mixed-motives, resource management, and interconnectedness easley2010networks; jensen2022complexity; thurner2018introduction, and produce networks that are dynamic, directed, weighted, and signed. Because these dynamics are encoded within the Junior High Game (JHG) (skaggs2024fostering), we use this strategic network game to study emergent behavior.
@@
A fundamental challenge in modeling human decision-making in the JHG lies in determining how to parameterize and represent behavior (rahwan2019machine). In this paper, we consider models that make different assumptions about human decision-making. One approach, behavior matching (axelrod1984evolution), assumes that agents imitate the observed actions of others, while another approach, community-aware behavior (skaggs2024fostering), considers broader structural and social cues within the network. We also examine how modeling different statistical moments (the mean verses the distribution of behavior) impact model performance. Together, this produces four different algorithms for modeling human behavior in the JHG. As a starting point to understanding how well these algorithms model human behavior, we compare and contrast their behavior in small-scale societies.
[…]
@@
Figure 1. Example scenario illustrating various JHG dynamics. (a) Popularity of the players over time. (b-c) Token allocations in rounds 1-4 and 5-6, respectively. Red dashed arrows indicate tokens used to attack while black solid arrows indicate giving.
We evaluate the ability of these modeling algorithms to produce human behavior in the JHG in two ways. First, we compare and contrast the population dynamics of agent societies with those of human groups. Results show that societies of one of these agents, called hCABs, have population dynamics that mirror (with some differences) human populations. Second, we conduct a user study to identify the extent to which individual hCABs exhibit human-like tendencies. Results of this study show that human participants have difficulty distinguishing hCAB agents from other human players. These findings not only inform the development of models that can simulate human behavior in the JHG, but also offer perspectives on the mechanisms that drive the dynamics of human networks.
2. A strategic network game
Human networks are punctuated by power asymmetry, mixed motives, resource management, and interconnectedness easley2010networks; jensen2022complexity; thurner2018introduction, which (along with other properties) combine to produce emergent properties of human networks. These characteristics are modeled in the Junior High Game (JHG) skaggs2024fostering, a strategic game that induces a dynamic (time varying), directed, weighted, and signed network.
We paraphrase skaggs2024fostering to review this game: The JHG is played by a set of players $I$ over a series of rounds in which each player $j\in I$ seeks to become popular. Initially, every player is assigned a popularity, which changes over time based on interactions, represented by token allocations, among the players. In every round, each player $j$ allocates $N$ tokens, each of which it can keep, give to another player, or attack another player. Keeping tokens positively impacts $j$’s subsequent popularity, while tokens $j$ gives to player $i$ positively impact $i$’s subsequent popularity (but do nothing for $j$). Finally, when $j$ attacks $i$, it negatively impacts $i$’s subsequent popularity while positively impacting $j$’s. After all players allocate their tokens in a round, the players’ popularities are updated based on the token allocations. A new round then begins.
The amount that a token transaction impacts subsequent popularity depends on the popularity of the player allocating the token. For example, receiving a token from a more popular player results in a higher increase in popularity than receiving a token from a less popular individual. Thus, players with higher popularity can have higher impact than those with lower popularity.
@@ 2. A strategic network game
To increase understanding of JHG dynamics, consider the example depicted in Figure 1. First, when players exchange (give) tokens with each other, they rise in popularity (e.g., players p2 and p3). Second, when a player gives tokens without receiving as much in return, they tend to decrease in popularity (e.g., p4). Third, when a player keeps its token, it maintains much of its popularity, but does not increase in popularity in the absence of other interactions (e.g., p1 in rounds 1-4). Fourth, attacking those not keeping tokens results in the attacker becoming more popular and the victim becoming less popularity (e.g., p5 attacking p7 and p8 in rounds 1-4). When other players do not keep or retaliate, attacking is the fastest way to gain popularity. It also increases the attacker’s relative popularity in relation to the victim. Fifth, keeping tokens blocks attacks. For example, p6’s attacks on p1 in rounds 1-4 are blocked because p1 is keeping its tokens–p1 experiences no loss, while p6 quickly loses popularity due to not gaining from its token allocations. Finally, attacks that are stronger than the defense (strength is measured as tokens allocated multiplied by player popularity) result in inflicted damage. For example, p5’s and p6’s combined attack on p1 in rounds 5-6 is substantially stronger than p1’s defense. Thus, p1 loses popularity and p5 and p6 profit somewhat.
@@
While this contrived example deals primarily with pairwise interactions for illustrative purposes, these dynamics can give rise to complex group behavior and emergent social norms. Frequently, alliances form and dissolve as players join together to help each other, punish those that violate norms, mitigate powerful enemies, and exploit others.
The complex strategic nature of the JHG gives rise to questions about how humans play it. Unlike many social dilemmas, players in the JHG must go beyond deciding whether to cooperate or defect. They must decide with whom they should cooperate (give tokens), when to create new relationships, when to dissolve friendships, when to defend one’s self (by keeping tokens), and who they should attack (take tokens) and when. In the next section, we describe methods that attempt to encode human behavior in the JHG.
3. Modeling human behavior in the JHG
@@
In this section, we introduce four algorithms that learn models of human behavior in the JHG by observing the play of experienced humans. These algorithms vary based on assumptions encoded in their parameter-based representations and methods for learning parameterizations. We discuss these design choices in turn.
3.1. Parameter-based models of behavior
We could potentially learn models of human behavior in the JHG using a variety of approaches. For instance, we could take a sort of tabula rasa approach, in which we feed large amounts of human data into a neural network. Unfortunately, obtaining large amounts of data of people playing the JHG (or some other strategic network game) is difficult. Our data set currently contains only a few hundred human player games, and it is unclear how to create human-like data synthetically. Thus, using a data hungry approach, such as a graph neural network, appears to be somewhat impractical. Additionally, we would like our model to be interpretable, meaning that it can inform us about the strategies used by human players. Black-box methods, such as graph neural networks, typically do not provide this desired interpretability.
As such, we consider the ability of parameter-based models to learn human behavior in the JHG. Such models are based on assumptions of how humans behave, with parameters that modulate behavior. Formally, let $m$ be the number of parameters in the model. Then, $\Theta=(\theta_{1},\theta_{2},\cdots,\theta_{m})$ is a parameterization of the model in which $\theta_{i}$ is the value of the $i^{\rm th}$ parameter.
In this paper, we consider two models: (1) Behavior Matching (or tit-for-tat; TFT) axelrod1984evolution and (2) Community-Aware Behavior (CAB) skaggs2024fostering.
@@ 3.1. Parameter-based models of behavior
Behavior Matching (TFT): In this model, a player reciprocates the actions that other players directed towards them in the previous round axelrod1984evolution. We consider this model due to its prevalence in prior work. TFT and its variants have repeatedly shown to be effective and robust in repeated prisoner’s dilemmas (e.g., axelrod1980effective; nowak1992tit; grant2013give; nowak1992evolutionary), a common domain for studying when and how humans and other agents cooperate with each other. Furthermore, TFT and its variants are frequently used to describe (successful) human behavior (e.g., axelrod1984evolution; grant2013give). As such, we consider whether human players use TFT-like strategies in the JHG, and, thus, whether a parameterized TFT algorithm can be tuned to successfully mimic human behavior.
@@
However, because a player’s resources are limited in the JHG, defining TFT in this domain is not straightforward, as a player may receive more or less in a round than it has the ability to reciprocate. In such cases, an agent must determine what to do with extra tokens (when receiving too much) or who to exclude (when not receiving enough). We have defined a TFT algorithm for the JHG whose behavior is tuned with seven parameters that determine how the agent allocates tokens in the first round, what it seeks to match (token allocations or influence), and how it deals with excess and shortfalls. Details are given in the supplementary material (SM-2).
Community-Aware Behavior (CAB): The second parameter-based model we consider is the CAB (Community Aware Behavior) algorithm (skaggs2024fostering). While using behavior matching to model humans assumes that humans cooperate and attack each other due to reciprocation, using the CAB algorithm assumes that humans help and attack each other based on group dynamics moffett2019human. Players in the same group support (give to) each other, work together to exploit other groups, and defend each other against the attacks of outsiders. CAB agents follow a parameterized algorithm for how humans form, join, and modify these groups over time.
@@
A CAB agent determines token allocations in a round using a three-step process. First, it assigns players to groups based on past token allocations. Second, it chooses a (potentially new) group it would like to form, support, or join. Finally, it selects token allocations designed to build up its selected group, establish itself in the group, and deal with out-of-group threats. CAB behavior is defined by thirty parameters which determine how it performs each of these steps. We use the CAB implementation provided by skaggs2024fostering.
@@
Assuming that human behavior in the JHG can be modeled using either of these parameterized algorithms (TFT or CAB) is most certainly incorrect at some level. However, identifying the extent to which these models can and cannot approximate human behavior can shed valuable insights into human networks and how human behavior can be effectively simulated in them.
3.2. Learning Parameterizations
@@
To determine how well TFT and CAB parameter models can model human behavior in the JHG, we learn parameterizations of each model from human data using two different methods: Particle Swarm Optimization (PSO) kennedy1995particle and Evolutionary Population Distribution Modeling (EPDM), a methodology proposed in this paper. Both algorithms take as input a parameter-based model $\mathcal{M}$ and output one or more parameterizations of that model.
@@
Particle Swarm Optimization (PSO): In PSO, a set of particles (i.e., parameterizations) are first randomly generated. Over time, these particles are adapted in a way that both explores the parameter space and converges toward the best parameterization, or the parameterization that minimizing the error, over all data samples, between the token allocations made by human players in the data and the token allocations made by the model (given the parameterization specified by the particle) in the same situations. PSO produces a parameterization designed to describe mean human behavior represented in the data.
@@
We experimented with two error (or distance) functions: (1) mean squared error and (2) a custom built function which measures how well a token allocation profile matches a target profile across a variety of properties. Because the second function outperformed the first in all cases, we used it in our experiments. This function scores the degree to which token allocation profile $\mathbf{b}$ matches the target profile $\mathbf{a}$ as follows:
${\mathcal{S}}(\mathbf{a},\mathbf{b})={\mathcal{S}}^{+}(\mathbf{a},\mathbf{b})+{\mathcal{S}}^{-}(\mathbf{a},\mathbf{b})+{\mathcal{S}}^{\rm keep}(\mathbf{a},\mathbf{b})-{\mathcal{P}}(\mathbf{a},\mathbf{b}).$ (1)
Here, ${\mathcal{S}}^{+}(\mathbf{a},\mathbf{b})$ scores the degree to which $\mathbf{b}$ (a) has positive allocations to the same number of players as $\mathbf{a}$, (b) has positive allocations to the same players as $\mathbf{a}$, (c) matches the total number of tokens allocated for giving as $\mathbf{a}$, and (d) gives the same number of tokens to each player as $\mathbf{a}$. ${\mathcal{S}}^{-}(\mathbf{a},\mathbf{b})$ scores the degree to which $\mathbf{b}$ (a) has a negative allocation if $\mathbf{a}$ has a negative allocation, (b) has the same number of tokens used for taking as $\mathbf{a}$, and (c) takes tokens from the same players as $\mathbf{a}$. ${\mathcal{S}}^{\rm keep}(\mathbf{a},\mathbf{b})$ compares the number of tokens kept in allocations $\mathbf{a}$ and $\mathbf{b}$. Finally, ${\mathcal{P}}(\mathbf{a},\mathbf{b})$ is a penalty term when allocation $\mathbf{b}$ gives to some player $i$ but allocation $\mathbf{a}$ takes from player $i$ or vice versa. Details related to ${\mathcal{S}}(\mathbf{a},\mathbf{b})$ are given in SM-4.
@@
Evolutionary Population Distribution Modeling (EPDM): Because mean human behavior, when followed by all individuals, may not produce human-like population dynamics or realistic individual behavior, we introduce a second method, called EPDM, for modeling the distribution of human behavior. The EPDM algorithm, outlined in Algorithm 1, receives as input a parameter-based behavior model $\mathcal{M}$ (we consider TFT and CAB in this paper) and a set of player games $\Gamma$ (from the training data). Each player game $\gamma\in\Gamma$ specifies the token allocations made by a human player in each round $t$ of a game, along with the game state available to the player at the beginning of that round. This game state consists of a vector ${\mathcal{P}}(t)$ specifying the current popularity of each player, a matrix ${\mathcal{I}}(t)$ specifying the influence that the players have on each other, and a matrix $X(t)$ specifying the token allocations made by all players in that round. The algorithm outputs a set of $N$ parameterizations of $\mathcal{M}$ that best estimate the distribution of player strategies in the player games $\Gamma$. We used $N=100$.
Algorithm 1 EPDM – Computes NN parameterizations to estimate a distribution of strategies
@@
EPDM learns this distribution of parameterizations using genetic evolution. Initially, a set of $N$ parameterizations (denoted $\Pi(0)$) are randomly generated. Over a series of $G$ generations, the pool of parameterizations is evolved to better estimate the distribution of strategies used by human players in the supplied player games. Each generation $g$ proceeds in three steps. In step 1, the set of player games for which parameterization $\pi\in\Pi(g)$ is a top performer is computed. This set, denote $\Gamma^{\pi}(g)$, is determined as follows. Let $\mathbf{a}^{\gamma}(t)$ denote the token allocations made in player game $\gamma\in\Gamma$ in round $t$. Also, let $\mathbf{a}^{\pi}(t)$ be the allocations that parameterization $\pi$ would make in round $t$ given the game history (prior to round $t$). Then, the average error of $\pi$ in player game $\gamma$ is
$\delta^{\pi}_{\gamma}=\frac{1}{T_{\gamma}}\sum_{t=1}^{T_{\gamma}}dist(\mathbf{a}^{\gamma}(t),\mathbf{a}^{\pi}(t)),$ (2)
@@
where $T_{\gamma}$ is the number of rounds in player game $\gamma$ and $dist(\mathbf{a},\mathbf{b})$ is the error function specifying the distance between allocations $\mathbf{a}$ and $\mathbf{b}$ (we use ${\mathcal{S}}(\mathbf{a},\mathbf{b})$; Eq. 1). $\pi$ is a top-performing parameterization for player game $\gamma$ if $\delta^{\pi}_{\gamma}-\min_{\pi^{\prime}\in\Pi}\delta^{\pi^{\prime}}_{\gamma}<\varepsilon$. We use $\varepsilon=0.01$.